OpenStudy (anonymous):

how do i solve the inequality -5x + 8 less than or greater to x - 4

OpenStudy (anonymous):

\[\LARGE -5x+8\leq x-4\] \[\LARGE -5x-x\leq -4-8\] only when you divide\multiply by a negative number , you swap the sign... otherwise it's just like an equation...

OpenStudy (anonymous):

\[\LARGE -6x\leq -12\] \[\LARGE 6x\geq 12\] now divide both sides by 6 and that's it..

OpenStudy (anonymous):

where did 6 come from